A man increased his speed by 5m/s after 30s,if his initial velocity was 20m/s.the average distance covered if his speed was at uniform acceleration would be?


1
Expert's answer
2020-10-12T07:45:02-0400

The distance covered by uniform acceleration is:


"d = v_0t + \dfrac{at^2}{2}"

where "v_0 = 20m\/s" is the initial speed, "t = 30s" is the time of motion, "a" is the uniform acceleration.

By definition, the acceleration is:


"a = \dfrac{\Delta v}{t}"

where "\Delta v = 5m/s" is the increase in the speed in time "t". Thus, obtain:

"d = v_0t + \dfrac{\Delta vt^2}{2t} = v_0t + \dfrac{\Delta vt}{2} = \left(v_0 + \dfrac{\Delta v}{2} \right)t"

In numbers:


"d = \left(v_0 + \dfrac{\Delta v}{2} \right)t = \left(20 + \dfrac{5}{2} \right)\cdot 30 = 675m"

Answer. 675 m.